OpenStudy (anonymous):

A baseball player is stealing second base when he goes into a slide, which gives him an acceleration of −2.5 m/s2. If he had an initial speed of 7 m/s, how far does he slide before coming to a complete stop?

OpenStudy (irishboy123):

use \[v^2 = u^2 +2ax\]

OpenStudy (shamim):

Given v=0 u=7m/s a=-2.5m/s^2 x=?